OpenStudy (anonymous):

Could somebody please help me with the following problem? Solve the equation tan(theta) = -2sin(theta) Also, isn't there an equation utility on this site? Thank you!

OpenStudy (anonymous):

Sorry, it also needs to be within the interval (0\[0\le \theta \le2\]

OpenStudy (anonymous):

\[\tan \theta=-2 \sin \theta\\ \frac{\sin \theta}{\cos \theta}=-2 \sin \theta \ \ \ \ \ \ \cos \theta \neq0\\ \sin \theta=-2 \sin \theta \cos \theta\\ \sin \theta+2 \sin \theta \cos \theta=0\\ \sin \theta(1+2 \ \cos \theta)=0 \\then \ \ u \ \ have \ \ sin \theta=0 \ \ or \ \ \ cos \theta=-\frac{1}{2}\]

OpenStudy (anonymous):

Thank you very much for the help, I really appreciate it. Also, I asked an additional question, "What is the solution within the interval [0,2\pi]

OpenStudy (anonymous):

well for sin theta=0 theta=0,pi,2pi and for cos theta =-1/2 theta=2pi/3,4pi/3

OpenStudy (anonymous):

Yeah, I was working right along with you and came up with same solution after working the problem approximately 300x's! =). Thank you sir! I really appreciate it!

OpenStudy (anonymous):

your welcome